OpenStudy (anonymous):

solve for T c=3t+1/t use the distributive property

OpenStudy (anonymous):

first distribute

OpenStudy (anonymous):

i dont know how :(

OpenStudy (anonymous):

\[\frac{ 3t+1 }{ t }=\frac{ 3t }{ t }+\frac{ 1 }{ t }\]

OpenStudy (anonymous):

would the answer be: t=1/c-3?

OpenStudy (anonymous):

yeah right :)